Grade 11Algebra

4 mothers each having one child walk for an interview.in how many ways this interview can be scheduled such that no child is interviewed before his mother.

bhavika
9 Years agoGrade 11

1 Answer

Saurabh Koranglekar
7 Years ago

To determine how many ways the interview can be scheduled with the condition that no child is interviewed before their mother, we can use a combinatorial approach. Let's break it down step by step.

The Setup

Imagine we have four mothers, each with one child: Mother 1 (M1) and Child 1 (C1), Mother 2 (M2) and Child 2 (C2), Mother 3 (M3) and Child 3 (C3), and Mother 4 (M4) and Child 4 (C4). We need to schedule interviews such that each child is interviewed only after their respective mother.

Understanding the Total Arrangements

If there were no restrictions, we could arrange the eight individuals (four mothers and four children) in any order. The total number of arrangements for eight distinct individuals is calculated using the factorial of the number of individuals:

Total arrangements without restrictions: 8! = 40,320

Applying the Condition

Now, we need to consider the requirement that each mother must be interviewed before her child. For any pair (M1, C1), (M2, C2), (M3, C3), and (M4, C4), there are exactly two arrangements: either the mother comes first or the child comes first. Since we want only the arrangement where the mother is interviewed before the child, we can think of this as a restriction on half of the arrangements.

Calculating Valid Arrangements

For each pair of mother and child, we have only one valid arrangement out of the two total arrangements. Since there are four pairs, we can apply this logic to our total arrangements:

Valid arrangements: Total arrangements divided by the number of arrangements for each pair = 8! / 2^4

Calculating this gives:

  • 8! = 40,320
  • 2^4 = 16
  • Valid arrangements = 40,320 / 16 = 2,520

Final Result

Thus, the total number of ways to schedule the interviews such that no child is interviewed before their mother is 2,520.

This combinatorial approach not only counts the arrangements but also respects the conditions imposed by the problem, allowing us to arrive at a precise and logical conclusion.